OpenStudy (anonymous):

derivative of e^((x^2) cos x)

OpenStudy (anonymous):

\[\frac{d}{dx}e^{x^2\cos x}=e^{x^2\cos x}\frac{d}{dx}(x^2\cos x)=e^{x^2\cos x}\left(\cos x\frac{d}{dx}[x^2]+x^2\frac{d}{dx}[\cos x]\right)=\cdots\]

OpenStudy (anonymous):

i got e^(x^2) cos x . 2x cos x .- sin x x^2 but wolfram alfa says somthing else

OpenStudy (anonymous):

Looks like you got\[\frac{d}{dx}e^{x^2\cos x}=\left(2x\cos x-x^2\sin x\right)e^{x^2\cos x}\] which is right. WA probably just factored something out, or tried an identity.